OpenStudy (anonymous):

solve the equation |k-10|=3

hartnn (hartnn):

when |a|=b a=b or a = -b does this help ?

OpenStudy (anonymous):

Yes, it should! I'd like to complete the solution, though. So, here, we have k - 10 = -3 or k - 10 = 3, so k = 7 or k = 13.

hartnn (hartnn):

let her do it

OpenStudy (anonymous):

how do i know what a and b is

hartnn (hartnn):

that is a general form. from that you get k-10 = 3 or k-10=-3 now solve both to get 2 values of k

OpenStudy (anonymous):

ok i get it now thank you!

hartnn (hartnn):

welcome ^_^